Danny Metzger's parents invested $1900 when he was born. This money is to be used for Danny's college education and is to be withdrawn in four equal annual payments beginning when Danny is age 19. Find the amount that will be available each year, if money is worth 7%, compounded annually. (Round your answer to the nearest cent.)
Answer #1

The formula for maturity value (F) of an initial deposit (P), after n years, where interest rate is r % and interest is compounded t times per year is F = P(1+r/100t)nt.

Here, P = $ 1900, r = 7/100 = 0.07 and n = 19. Therefore, F = 1900(1+7/100)19 = 1900* 3.616527535 = $ 6871.40 (on rounding off to the nearest cent).

The annuity payment formula is P = r(PV)/[ 1-(1+r)-n] where PV is the present value, P is the periodic payment, r is the rate per period and n is the number of periods.

Here, PV = $ 6871.40, r = 7/100 = 0.07 and n = 4.

Hence, P = (0.07* 6871.40)/[ 1-(1.07)-4] = (0.07* 6871.40)/0.237104788 = $ 2028.63 (on rounding off to the nearest cent).

Thus, the amount that will be available each year is $ 2028.63.